# Continuous-variable quantum key distribution with noisy squeezed states

### v1

@@ 5 Anti-squeezing noise in fluctuating channels
While in the previous sections we assumed the channel transmittance to be stable, which typically corresponds to fiber-type channels, in the free-space realizations it is usually not the case and the atmospheric channel transmittance fluctuates due to the turbulence effects, we further refer to such transmittance fluctuations as to the channel fading. Although the changes of transmittance are slower than the clock rate of the protocol, such attenuation fluctuations during data accumulation can significantly impede achievable secure key rate [20, 21]. The overall state shared between trusted parties after the fading channel is generally non-Gaussian and is described by the Wigner function, which is the weighed convex sum of Wigner functions that characterize shared state when Bob’s subsystem was transmitted through a channel with fixed transmittance $T_{i}$ that occurs with probability $\tau_{i}$, with $\sum^{M}_{i=1}\tau_{i}=1$ [40]. While the first moments remain zero, the second moments $\gamma_{AB}$ of the resulting state are hence the convex combination with transmittance $T_{i}$ and transmittance coefficient $\sqrt{T_{i}}$ of all $i\in[1,M]$. Extremality of Gaussian states [41] and subsequent optimality of Gaussian collective attacks [32, 33] allows us to gaussify the overall mixed state, i.e. base further security analysis on the covariance matrix $\gamma_{AB}$, which is now evidently a function of $\langle T\rangle$ and $\langle\sqrt{T}\rangle$, without underestimating the eavesdropper. In the current section we consider the homodyne-based protocol as the best performing in the asymptotic regime, and allowing to reach longer distances in finite-size regime. Bob’s subsystem after the fading channel is described by
@@
A fading channel can be equivalently represented as a channel with fixed transmittance $\langle\sqrt{T}\rangle^{2}$ and fading-related variance-dependent excess noise $\varepsilon_{x}^{f}=Var(\sqrt{T})(V+V_{x}-1)$ and $\varepsilon_{p}^{f}=Var(\sqrt{T})(1/V+V_{p}+\Delta V-1)$ in $x$- and $p$-quadratures respectively, contrary to symmetrical fading-related noise in the coherent-state protocol [20]. The noise related to fading is to be assumed untrusted, as the transmittance fluctuations can be fully controlled by Eve. This represents another strongly negative impact of AS noise in the fluctuating channels, even if the noisy source is initially trusted. Such noise necessitates the characterization of transmittance coefficient variance $Var(\sqrt{T})=\langle T\rangle-\langle\sqrt{T}\rangle^{2}$ and consequent adaptation of modulation $V_{x(p)}$ and squeezing variance $V$ for optimal secure key rate. For weak atmospheric turbulence one can expect $Var(\sqrt{T})\leq 0.01$, and $Var(\sqrt{T})>0.04$ for strong turbulence [21]. Mutual information (21) reduces to (2) when fluctuations of transmittance are negligible, i.e. $M=1$ and $\tau_{i=1}=1$ or $Var(\sqrt{T})=0$. The upper bound on Eve’s accessible information $\chi_{AE}$ or $\chi_{BE}$ is evaluated based on $\gamma_{AB}(\langle T\rangle,Var(\sqrt{T}))$ as well. For further details regarding the modelling and security analysis of atmospheric channels in CV QKD see [20].
@@
6 Conclusion
In this work, we perform an exhaustive investigation into the role of the noise in squeezed signal states in CV QKD, considering collective attacks in the asymptotic as well as finite-size regimes, and incorporating free-space channel fading. We show how the noise in the signal squeezing can unavoidably appear as the result of optical losses already in the state generation. The impact of this noise however greatly depends on whether it can be assumed trusted, i.e., originating from the trusted source and not controlled by an eavesdropper, or not. In the case, when this noise is trusted, we firmly assume that it is contained in the AS quadrature. Such trusted noise can even slightly reduce the Holevo bound, leading to analytical convergence to individual attacks in the limit of strong noise. The trusted AS noise can also be helpful in DR scenario as a counter-measure against the channel noise. In the finite-size regime, however, already the trusted AS noise can limit the efficiency of parameter estimation and requires optimization of the protocol and the estimation strategies. Considering biased homodyne detection and unbalanced heterodyne detection as the possible strategies on the receiver side and taking into account the possibility to announce the outcomes of the AS quadrature measurements, we identify the optimal protocol implementations. Our findings show that biased alternation between homodyne measurements of squeezed and AS quadrature of signal field, followed by disclosure of the latter, allows to extend the secure distance. However, simultaneous measurement of both quadratures can still be advantageous in certain regimes, especially for shorter data accumulation times, resulting in shorter data block sizes. Finally, we consider the fading channels with fluctuating transmittance, which is typical for free-space atmospheric channels, and show that even trusted AS noise transforms to the untrusted AS noise due to fading. This emphasizes the importance of the state purity for squeezed-state CV QKD realizations in free-space channels already with the trusted sources. Our findings present the methods for protocol optimizations, including measurement and estimation techniques, as well as suggest the experimental efforts towards the signal state purity for efficient squeezed-state CV QKD. The obtained results pave the way towards practical and efficient realization of the squeezed-state QKD, they can be directly applied in point-to-point protocols as well as in QKD-based communication networks, where great diversity of implementation conditions and combination of various protocols and link types is anticipated.
@@ Acknowledgements
A. O. and V. C. U. acknowledge the project 8C22002 (CVStar) of MEYS of Czech Republic, which has received funding from the European Union’s Horizon 2020 research and innovation framework programme under grant agreement No. 731473 and 101017733, and the project No. 21-44815L of the Czech Science Foundation. A. O. acknowledges the project IGA-PrF-2024-008 of Palacky University Olomouc. I. D. acknowledges the project 22-28254O of the Czech Science Foundation. V. U. acknowledges the European Union’s Horizon Europe research and innovation programme under the project ”Quantum Security Networks Partnership” (QSNP, grant agreement No. 101114043).
Appendix A Squeezed-state noise allocation
@@
To gain insights into the impact of arbitrary noise distributions, we perform numerical investigations focusing on impure squeezed state obtained due to attenuation of pure squeezed state Fig. 1. Fig. 10 displays the Holevo bound versus excess squeezed quadrature variance $\epsilon_{x}$ for various transmittance values and degrees of squeezing, with the excess noise in the (AS) quadrature being $\epsilon_{p}=1+\eta(1/V_{s}-1)-1/(1+\eta[V_{s}-1]-\epsilon_{x})$. The plots provide intriguing insights into the consequences of employing an impure squeezed state. One notable observation is that, under worst-case considerations, increased impurity does not necessarily translate to reduced security. This phenomenon stems from the fact that a higher impurity in a noisy squeezed state does not inherently result in higher AS noise. To illustrate, consider the attenuation of a highly squeezed state ($V_{s}$ = 0.1 SNU) with beamsplitter transmittance values $\eta$ = 0.5 and $\eta$ = 0.7. Despite the state with higher losses being less pure, the maximum AS excess noise for $\eta$ = 0.5 is less than that for $\eta$ = 0.7.
@@
Figure 11: (a) shows optimal $\epsilon_{x}$ to obtain minimal key rate Eq. 23, for different noisy squeezed state with squeezed quadrature variance $\tilde{V}_{s}$, and AS quadrature variance $\tilde{V}_{AS}=10$ SNU (black) and $\tilde{V}_{AS}=5$ SNU (blue). (b) shows key rates for DR, when optimal noise distribution from (a) is used (dashed) and when all the noise is considered to be in AS quadrature (solid).
Appendix B Parameter estimation

### v2

@@ 4 Anti-squeezing noise in fluctuating channels
While in the previous sections we assumed the channel transmittance to be stable, which typically corresponds to fiber-type channels, in the free-space realizations it is usually not the case and the atmospheric channel transmittance fluctuates due to the turbulence effects, we further refer to such transmittance fluctuations as to the channel fading. Although the changes of transmittance are slower than the clock rate of the protocol, such attenuation fluctuations during data accumulation can significantly impede achievable secure key rate [27, 28]. The overall state shared between trusted parties after the fading channel is generally non-Gaussian and is described by the Wigner function,
which is the weighed convex sum of Wigner functions that characterize shared state when Bob’s subsystem was transmitted through a channel with fixed transmittance $T_{i}$ that occurs with probability $\tau_{i}$, with $\sum^{M}_{i=1}\tau_{i}=1$ [48]. While the first moments remain zero, the second moments $\gamma_{AB}$ of the resulting state are hence the convex combination with transmittance $T_{i}$ and transmittance coefficient $\sqrt{T_{i}}$ of all $i\in[1,M]$. Extremality of Gaussian states [49] and subsequent optimality of Gaussian collective attacks [40, 41] allows us to gaussify the overall mixed state, i.e. base further security analysis on the covariance matrix $\gamma_{AB}$, which is now evidently a function of $\langle T\rangle$ and $\langle\sqrt{T}\rangle$, without underestimating the eavesdropper. In the current section we consider the homodyne-based protocol as the best performing in the asymptotic regime, and allowing to reach longer distances in finite-size regime. Bob’s subsystem after the fading channel is described by
@@
A fading channel can be equivalently represented as a channel with fixed transmittance $\langle\sqrt{T}\rangle^{2}$ and fading-related variance-dependent excess noise $\varepsilon_{x}^{f}=Var(\sqrt{T})(V+V_{x}-1)$ and $\varepsilon_{p}^{f}=Var(\sqrt{T})(1/V+V_{p}+\Delta V-1)$ in $x$- and $p$-quadratures respectively, contrary to symmetrical fading-related noise in the coherent-state protocol [27]. The noise related to fading is to be assumed untrusted, as the transmittance fluctuations can be fully controlled by Eve. This represents another strongly negative impact of AS noise in the fluctuating channels, even if the noisy source is initially trusted. Such noise necessitates the characterization of transmittance coefficient variance $Var(\sqrt{T})=\langle T\rangle-\langle\sqrt{T}\rangle^{2}$ and consequent adaptation of modulation $V_{x(p)}$ and squeezing variance $V$ for optimal secure key rate. We compare all noises involved in the protocol and highlight the differences between them, as well as methods for their estimation and their influence on the security in Tab. 1. For weak atmospheric turbulence one can expect $Var(\sqrt{T})\leq 0.01$, and $Var(\sqrt{T})>0.04$ for strong turbulence [28]. Mutual information (21) reduces to (2) when fluctuations of transmittance are negligible, i.e. $M=1$ and $\tau_{i=1}=1$ or $Var(\sqrt{T})=0$. The upper bound on Eve’s accessible information $\chi_{AE}$ or $\chi_{BE}$ is evaluated based on $\gamma_{AB}(\langle T\rangle,Var(\sqrt{T}))$ as well. For further details regarding the modelling and security analysis of atmospheric channels in CV QKD see [27].
@@
Noise Excess $\epsilon$ Fluctuation $\varepsilon^{f}$ Anti-squeezing $\Delta V$ Detector noise $\nu_{el}$
Origin
Untrusted
channel
Gaussification
Cavity loss,
preparation
imperfections
Electronic noise
Estimation
method
Data disclosure
Separate measurement
and/or
directly from data
Data disclosure
and/or
during calibration
Separate
measurement
Influence See Fig. 6 and 5 See Fig. 9 and 8 See Fig. 7 See Fig. 3
Trust No No Up to users Yes
Table 1: Comparison of different noises involved in the squeezed-state CV QKD protocol.
@@
5 Summary and conclusions
In this work, we perform a detailed investigation into the role of the excess noise in squeezed states in CV QKD, studying security against collective attacks in the asymptotic as well as finite-size regimes, and considering free-space channel fading. In the widely adopted trust assumption on the noisy squeezing, we show the robustness of the protocol against squeezing excess noise in the asymptotic regime of infinitely many data points, first for the purely attenuating channels, where strong trusted squeezing noise allows to obtain compact analytical expressions for the Holevo bound, as well as in the presence of channel noise. Furthermore, we largely confirm the stability of protocol against trusted noisy squeezing in the finite-size regime, where we consider two main strategies for the parameter estimation:
  1. 1.
    Biased homodyne detection, where Bob alternates between measuring the squeezed and AS quadratures, and
  2. 2.
    Imbalanced heterodyne detection, where Bob performs double homodyne detection with a beam splitter transmission skewed towards measuring the squeezed quadrature.
We find that the optimal method depends on the levels of squeezing, block size, and the amount of excess noise. Specifically, for moderately squeezed states $V=0.5$ SNU with low excess noise $\epsilon<0.1$ SNU, biased homodyne measurements enable key generation over the greatest distances. For all other scenarios, imbalanced heterodyne measurement proves to be more effective in achieving longer secure communication distances. In the case, when the noisy squeezing is untrusted, it introduces a security bound on the protocol realization, which is further enforced by the finite-size effects. In the case of untrusted noise, the purity of the signal squeezed states becomes crucial already in the asymptotic regime, especially in the reverse reconciliation scenario, aimed at long-distance realizations of CV QKD. Finally, we consider the fading channels with fluctuating transmittance, which is typical for free-space atmospheric channels, and show that even trusted squeezing excess noise transforms to the untrusted channel noise due to fading. This emphasizes the importance of the state purity for squeezed-state CV QKD realizations in free-space channels already with the trusted sources.
Our findings present the methods for protocol optimizations, including measurement and estimation techniques, targeting the practicality and efficiency of the protocol realization. The obtained results pave the way towards optimized experimental implementation of the squeezed-state QKD protocol. Our theory shows several opportunities for experimental improvements. Firstly, increasing the signal state purity can translate into a higher performance of the QKD protocol. Secondly, stability of the squeezed light generation is essential for a relivable protocol operation, as security assessment is contingent on precise characterization of the signal state throughout whole key generation process. Lastly, any phase misalignment at the receiver side can diminish the benefits of the squeezed state. Albeit, the latter can be mitigated by employing machine learning techniques [19]. We believe that this work will enable secure quantum communication based on squeezed-state of light
in point-to-point protocols as well as in QKD-based communication networks, where great diversity of implementation conditions and combination of various protocols and link types is anticipated.
@@ Acknowledgements
A. O. and V. C. U. acknowledge the project 8C22002 (CVStar) and R.F. acknoledges the project 8C22001 (SPARQL) of MEYS of Czech Republic, which have received funding from the European Union’s Horizon 2020 research and innovation framework programme under grant agreement No. 731473 and 101017733. V.C.U. also acknolwedges the project No. 21-44815L of the Czech Science Foundation. A. O. acknowledges the project IGA-PrF-2024-008 of Palacky University Olomouc. I. D. acknowledges the project 22-28254O of the Czech Science Foundation. V. U. acknowledges the European Union’s Horizon Europe research and innovation programme under the project ”Quantum Security Networks Partnership” (QSNP, grant agreement No. 101114043). R.F. and V.U. acknowledge the project CZ.02.01.01/00/22_008/0004649 (QUEENTEC) of MEYS of Czech Republic.
Appendix A Noisy squeezing due to optical losses
Squeezed states, typically generated using optical parametric oscillators [22], are unavoidably noisy already due to the optical losses in the cavity, phase noise, parasitic nonlinearities, absorptions, and inefficient collection of light at the cavity output. We model these effects in Gaussian approximation by propagating an effective pure squeezed state through an attenuation to reach the considered realistic squeezed state with an excess noise. A single-mode pure squeezed state, described by the squeezed and AS variances $V_{S},1/V_{S}$, undergoing losses can be modeled as coupling to a vacuum mode on a beamsplitter with transmittance $\eta$. The resulting state after the beamsplitter interaction [51] is described by the quadrature variances $\eta V_{S}+1-\eta,\eta/V_{S}+1-\eta$ and contains the fraction of vacuum noise (in shot-noise units, SNU), scaled by the beamsplitter reflectance $1-\eta$. The Gaussian purity of the state then becomes $\mu=\left[1+(\eta-\eta^{2})(1/V+V-2)\right]^{-1/2}$ and is minimized at $\eta=1/2$. When detected, without precise knowledge of coupling loss $\eta$ and initial squeezing $V_{S}$, one cannot pinpoint the exact distribution of added noise ($\epsilon_{x}$ and $\epsilon_{p}$) in each quadrature. In the context of CV QKD, when the actual $V_{S}$ and $\eta$ are not known, it can be assumed that the resulting state contains a pure squeezed state with variances $V,1/V$ and some excess noise $\epsilon_{x},\epsilon_{p}$ in either of the quadratures. The allocation of the noise once it is trusted (not controlled by Eve) does not influence the security of CV QKD. However, when the noise in the squeezed signal states is not trusted (controlled by Eve), its allocation to particular quadratures has direct security ramifications and we must adopt the most conservative stance. The latter typically implies admitting the largest quantitative amount of noise, i.e. treating all the noise as contained in the AS quadrature, see Appendix A for further details.
Appendix B Squeezed-state noise allocation
@@
To gain insights into the impact of arbitrary noise distributions, we perform numerical investigations focusing on impure squeezed state obtained due to attenuation of pure squeezed state Fig. 10. Fig. 11 displays the Holevo bound versus excess squeezed quadrature variance $\epsilon_{x}$ for various transmittance values and degrees of squeezing, with the excess noise in the (AS) quadrature being $\epsilon_{p}=1+\eta(1/V_{s}-1)-1/(1+\eta[V_{s}-1]-\epsilon_{x})$. The plots provide insights into the consequences of employing an impure squeezed state. One notable observation is that, under worst-case considerations, increased impurity does not necessarily translate to reduced security. This phenomenon stems from the fact that a higher impurity in a noisy squeezed state does not inherently result in higher AS noise. To illustrate, consider the attenuation of a highly squeezed state ($V_{s}$ = 0.1 SNU) with beamsplitter transmittance values $\eta$ = 0.5 and $\eta$ = 0.7. Despite the state with higher losses being less pure, the maximum AS excess noise for $\eta$ = 0.5 is less than that for $\eta$ = 0.7.
@@
Fig. 12(c) illustrates the impact of the squeezed state’s purity on the protocol’s security. It should be noted that, as previously explained and demonstrated, the security of the protocol hinges on the noise distribution causing the impurity. The plots display the key rates for the worst-case noise distribution corresponding to that purity.
@@
Figure 12: (a) shows optimal $\epsilon_{x}$ to obtain minimal key rate Eq. 23, for different noisy squeezed state with squeezed quadrature variance $\tilde{V}_{s}$, and AS quadrature variance $\tilde{V}_{AS}=10$ SNU (black) and $\tilde{V}_{AS}=5$ SNU (blue). (b) shows key rates for DR, when optimal noise distribution from (a) is used (dashed) and when all the noise is considered to be in AS quadrature (solid). (c) shows the key rates with a pessimistic noise distribution for RR (red) and DR (green), varying the purity of the squeezed state.
Appendix C Parameter estimation